OpenStudy (anonymous):

help pls what is the ph when 100ml of 0.1 m naoh is added to 150ml of 0.2m ch3cooh? pka for ch3cooh = 4.76

OpenStudy (anonymous):

Use the equation below. Figure out moles NaOH and CH3COOH and plug them in. \[pH = pKa + \log \left( \frac{ base }{ acid } \right) \]

OpenStudy (vincent-lyon.fr):

Have you written the reaction taking place during the titration?